### v2

@@ 8.5.1 Masking Percentage
8.5.2 Few-Shot Distance Metric
@@
9 Encrypted Traffic
@@
Despite these encryption properties, some studies have successfully classified encrypted network traffic from various applications [45, 46]. This classification success is largely due to the unique encryption signatures that different applications’ random generators leave, which, while not disclosing plaintext content, offer a form of application-specific pattern. The challenge, however, lies in malware detection, where access to plaintext information is crucial. If malware signatures can be detected in encrypted traffic, it would suggest that the encryption method fails to adequately mask the plaintext, thereby violating the principle stated in Equation 11. Our research involved applying two encryption methods to the packet data: the Advanced Encryption Standard (AES) and Fernet symmetric encryption [47, 48]. Fig. 7 represents the different pre-processing methods between both encryption methods.
@@ 10 Discussion
11 Conclusion
@@
This paper explored the application of a self-supervised transformer-based model with a downstream task and few-shot learning for malware detection and classification. The model was evaluated on the UNSW-NB15 and CIC-IOT23 datasets, focusing on the packets of UDP and TCP packets serving as inputs. Our method showed robust results on the classification of benign versus malicious packets when compared to SOA methods on both datasets. The few-shot results were also promising, showing the model’s ability to learn with limited amounts of labeled data. Using the packet bytes as input vectors, we were able to classify when different packets are either benign or malicious, as well as which types of attacks were present. While the payload bytes of packets are inherently different from the natural language structure, this study shows that the transformer-based model developed for natural language can be leveraged to capture and learn the intricate sequential patterns of the packet bytes.

### v3

@@ IX-A Masking Percentage
IX-B Few-Shot Distance Metric
@@
IX-C Sensitivity to Transformer Parameters
To investigate the sensitivity to transformer depth and embedding dimension, we conduct an ablation study to evaluate the sensitivity of our model to the depth of the transformer backbone and the dimensionality of its embeddings. Our original architecture is based on a 12-layer transformer with a hidden size of 768, consistent with standard base configurations. To simulate resource-constrained settings for SSL pre-training, we designed variants with reduced numbers of transformer layers and hidden sizes.
TABLE VIII: Transformer depth and dimension ablation on UNSW-NB15 for binary, multiclass, and FSL classification.
Depth / Dim Binary Multiclass FSL (5-shot)
Acc. F1 Acc. F1 Acc. Unseen Acc.
6 / 384 76.03 75.06 85.46 85.45 68.26 67.52
6 / 768 92.57 92.57 89.31 89.31 82.88 82.60
12 / 384 82.81 82.60 84.73 84.68 70.19 69.42
12 / 768 (baseline) 94.88 94.88 91.29 92.29 88.83 88.93
The rationale behind this design consists of two main reasons. First, the number of transformer layers in the model influences the representational capacity of the model, while also corresponding to the number of attention heads the model can apply. This enables deeper networks to capture longer-range dependencies and better refine attention patterns, although at the cost of increased computational overhead during pre-training. Second, the size of the embedding dimension determines how much information is retained in the latent space. Smaller dimensions may improve computational efficiency, but lose fine-grained and discriminative characteristics of the traffic bytes. By evaluating different configurations, we can better understand the relationships between model depth, representational power, and classification performance across binary, multiclass, and FSL scenarios.
Table 8 summarizes the results of this ablation on the UNSW-NB15 dataset, compared to baseline classification performance of each experiment. We report accuracy and F1 metrics across binary and multiclass experiments, and overall and unseen class accuracies for the FSL experiment based on the 5-shot baseline results, averaged over all three FSL experiments. Reducing transformer layer depth from 12 to 6 layers while retaining the 768-dimensional embedding yields the smallest degradation across all tasks. In contrast, narrowing the embedding to 384 dimensions across either depth substantially decreases performance, with the largest drops occurring in the binary and FSL classification tasks. The baseline models remains best overall, but among the reduced-capacity models, we can conclude that the 6/768 variant offers a favorable trade-off.
This occurs due to the smaller embedding dimension constraining the model’s capacity to encode discriminative features from the traffic data, especially under few-shot scenarios, where the model must adapt to the novel class that was not seen during pre-training. Consequently, this leads to the largest performance drop in the few-shot setting for average and unseen accuracies. Regarding the depth of the layers, we can conclude that most discriminative features are learned in the earlier layers of the model. However, the baseline performance of 12/768 still yields favorable results compared to the constrained variants.
X Encrypted Traffic
@@
Despite these encryption properties, some studies have successfully classified encrypted network traffic from various applications [46, 47]. This classification success is largely due to the unique encryption signatures that different applications’ random generators leave, which, while not disclosing plaintext content, offer a form of application-specific pattern. The challenge, however, lies in malware detection, where access to plaintext information is crucial. If malware signatures can be detected in encrypted traffic, it would suggest that the encryption method fails to adequately mask the plaintext, thereby violating the principle stated in Equation 11. Our research involved applying two encryption methods to the packet data: the Advanced Encryption Standard (AES) and Fernet symmetric encryption [48, 49]. Unlike our earlier work [21], which introduced text-encoding artifacts through character-level tokenization, this byte-level formulation avoids dataset-specific cues and provides a more rigorous basis for evaluating ciphertext learnability. Fig. 8 represents the different pre-processing methods between both encryption methods.
@@ XI-A Discussion
An additional observation is that the simple ML classifier consistently outperforms more complex learning architectures such as CNNs and LSTMs when leveraging the SSL embeddings, despite its relatively lightweight structure. This can be explained by the strength of the SSL-generated embeddings, which already encode local and global dependencies within packet data. Since these embeddings are highly discriminative, the ML classifier can leverage them directly without requiring additional temporal feature extraction. In contrast, the deeper models may introduce unnecessary complexity or risk overfitting on the SSL embeddings. This finding demonstrates that once robust embeddings are learned, a simple downstream classifier can lead to improved accuracy and faster inference, which is especially beneficial for real-time deployment.
@@
XI-B Limitations and Threats to Validity
Our study demonstrates promising results, but several limitations and threats to validity should be acknowledged. Pre-training with transformers is computationally expensive and depends heavily on dataset scale [50, 51]. Larger and more diverse datasets generally yield richer embeddings but require substantial compute and training time, which may limit accessibility for organizations without high-performance infrastructure. However, the growing availability of high-performance GPUs is making large-scale pre-training on datasets of benign and malicious behaviors increasingly feasible. In addition, because our SSL embeddings are fixed after pre-training, downstream tasks are constrained by the representations learned from the pre-training dataset. This can limit the model’s ability to continuously generalize to novel malware families over time. While our proposed FSL method partially mitigates this, future work could investigate incremental or online SSL training to refine embeddings in real time. With parameter-efficient fine-tuning methods such as LoRA [52]), this becomes increasingly practical, since only a small set of additive parameters must be updated rather than fully fine-tuning the entire model.
Furthermore, our approach assumes access to plaintext transport-layer payloads for DPI. In environments with widespread encryption, payload observability is reduced, and our model performs at chance when packet bytes are fully encrypted using Fernet or AES. We note earlier in the manuscript that DPI often presumes access to unencrypted payloads [31, 32], a condition that can still hold in certain enterprise networks, industrial control systems (ICS), and CAN applications. Future work will combine payload-agnostic statistics with our SSL embeddings and evaluate on further encrypted captures.
We trained and evaluated on two public corpora (UNSW-NB15 and CIC-IoT23) with deliberate class selection and de-duplication to balance classes; nonetheless, focusing on a limited set of attack families per corpus and sampling subsets can introduce bias in byte distributions and background traffic characteristics. While performance and few-shot transfer indicate promise within the chosen attacks across these two datasets, external validity remains limited to these settings. To strengthen generalization claims, future work will conduct cross-generalization capabilities of the proposed model, incorporate additional captures spanning network traffic collection pipelines, and evaluate domain-adaptation and deployability strategies for out-of-distribution real-time malware detection.
XII Conclusion and Future Research Directions
@@
This paper explored the application of a self-supervised transformer-based model with a downstream task and few-shot learning for malware detection and classification. The model was evaluated on the UNSW-NB15 and CIC-IOT23 datasets, focusing on the packets of UDP and TCP packets serving as inputs. Our method showed robust results on the classification of benign versus malicious packets when compared to SOA methods on both datasets. Furthermore, our results show that a simple ML classifier, when paired with SSL-generated embeddings, can outperform more complex CNN and LSTM models. This emphasizes that SSL pre-training results in discriminative embeddings, allowing lightweight models to achieve improved accuracy and efficiency compared to deeper architectures, which is advantageous for real-time intrusion detection. The few-shot results were also promising, demonstrating the model’s ability to learn with limited amounts of labeled data. Using the packet bytes as input vectors, we were able to classify when different packets are either benign or malicious, as well as which types of attacks were present. While the payload bytes of packets are inherently different from the natural language structure, this study shows that the transformer-based model developed for natural language can be leveraged to capture and learn the intricate sequential patterns of the packet bytes.
Despite these encouraging results, several challenges and opportunities remain. Although our evaluation focused on traditional enterprise network and IoT packets, it is crucial to extend out work to additional protocols across several domains. In particular, protocols associated with ICS and in-vehicle CAN present unique characteristics which may require additional pre-training. Furthermore, another promising direction is the design and implementation of new models based off our proposed transformer-based architecture that incorporates emerging paradigms such as continual learning, which would enable rapid adaptation to evolving traffic patterns while remaining suitable for deployment in resource-constrained environments. Lastly, building attribution models that leverage distinctive network packet signatures represents an important challenge, as such models would enhance interpretability by linking malicious behaviors to identifiable network traits. Overall, this work proposed a self-supervised transformer-based pre-training framework for malware detection and classification, demonstrating its versatility through multiple downstream tasks, including FSL.